OpenStudy (anonymous):

I have the answer can someone please check me... What is the derivative of p(x)=square root f(x)

OpenStudy (anonymous):

\[p(x)=\sqrt{f(x)}\]

OpenStudy (anonymous):

the answer that I got was\[\frac{ 1 }{ 2f ^{\frac{ 3 }{ 2 }}x ^{\frac{ 1 }{ 2 }} }\]

OpenStudy (anonymous):

f(x) is a function. The answer you want should be \[ \frac{d}{dx} \sqrt{f(x)} = \frac{1}{2\cdot \sqrt{f(x)}} \cdot f'(x) \]

OpenStudy (anonymous):

ok then how do I find p'(3)

OpenStudy (anonymous):

what is f(x)?